# R-PINN: Recovery-type a-posteriori estimator enhanced adaptive PINN

### v1

@@ 3 Recovery-type estimator enhanced PINN
In this section, we present an adaptive PINNs algorithm developed by our team, which utilizes the recovery-type a-posteriori error estimator [27, 28, 29], a widely used technique in adaptive finite element methods. The concept of adaptive finite element method, first introduced by I. Babuška et al. in 1978 [25], has since seen significant advancements and remains integral to practical scientific computations. This approach extends classical finite element methods by incorporating a-posteriori error estimation and automated mesh refinement to iteratively improve the accuracy of approximate solutions, which can achieve the target accuracy with reduced computational effort, offering distinct advantages over traditional finite element methods, particularly when addressing problems involving singular solutions.
The fundamental concept of the adaptive finite element method lies in the use of a-posteriori error estimator to assess the numerical solution error within each mesh element, thereby guiding the mesh refinement and coarsening process. This adaptive mesh strategy shares a notable similarity with adaptive PINNs, where the distribution of collocation points is adjusted based on the error distribution. In adaptive finite element methods, commonly employed a-posteriori error estimators can be categorized into two main types: the first is based on residual type [25, 36], where the error is estimated by computing the discrepancy between the solution and the residual of the governing equation; the second type, recovery type, proposed by the Zienkiewicz and Zhu [27, 28, 29], is a gradient reconstruction-based method that provides more accurate error estimates by reconstructing the solution gradients.
In existing adaptive PINNs algorithms, the equation operator is typically used as the default a-posteriori error estimator, which corresponds to the residual-based error estimator in adaptive finite element methods. Our proposed algorithm seeks to introduce the recovery-type error estimator into the adaptive PINNs framework, with the aim of enriching and extending the theoretical foundation of adaptive PINNs, ultimately providing a more comprehensive approach to error estimation and control.
@@ 3.1 Recovery-type a-posteriori estimator
In 1987, O.C. Zienkiewicz and J.Z. Zhu introduced a pioneering a-posteriori error estimator based on gradient recovery for finite element theory, commonly referred to as the Z-Z method [27] or the recovery-type method. This approach, notable for its computational simplicity, intuitive understanding, and straightforward implementation within finite element software, quickly gained widespread acceptance within the engineering community. Its rapid adoption represented a significant milestone, facilitating the transition from theoretical research to practical application in the domain of finite element a-posteriori error estimation.
In the context of a specific recovery-type a-posteriori error estimation framework, a typical problem in the field of elasticity is to be considered as an example, which is encapsulated within the following system of equations
$\bm{L}\bm{u}=\bm{S}^{T}\bm{DS}\bm{u}=\bm{f}\quad\text{in}\ \Omega,$ (3.1)
which is subject to suitable boundary conditions on the domain’s boundary $\partial\Omega$. Here, $\bm{u}$ represents the displacement vector, $\bm{L}$ is the differential operator intrinsic to the theory of elasticity, and the matrix operator $\bm{S}$ defines the strain tensor $\bm{\epsilon}$ by
$\bm{\epsilon}=\bm{S}\bm{u}.$ (3.2)
Furthermore, the elasticity matrix $\bm{D}$ is utilized to determine the stress tensor $\bm{\sigma}$, as delineated by the equation:
$\bm{\sigma}=\bm{D}\bm{\epsilon}.$ (3.3)
The finite element solution to the aforementioned elasticity problem is expressed in the following form
$\bm{u}_{h}=\bm{N}\bar{\bm{u}}_{h},$ (3.4)
where $\bar{\bm{u}}_{h}$ denotes the vector of nodal displacements, and $\bm{N}$ represents the set of approximate polynomial shape functions that interpolate the displacement field within the finite element framework Subsequently, the stress tensor $\bm{\sigma}_{h}$, is computed using the approximate displacement field $\bm{u}_{h}$ with
$\bm{\sigma}_{h}=\bm{DSu}_{h}.$ (3.5)
The recovery-type a-posteriori error estimator is characterized by focusing on estimating the numerical error by approximating the gradient recovery of the numerical solution $\bm{\sigma}^{*}$. Therefore, the gradient recovery technique plays a key role for the estimator. With the development of recovery-type a-posteriori error estimators, there have been many gradient recovery methods, the typical ones are the transformation method and nodal averaging method.
Fig. 2: The finite element method solution and the smoothened solution of stress for a one-dimensional problem.
  • 1.
    The transformation method
    For this method, it is assumed that the stress
    $\bm{\sigma}^{*}=\bm{N}\bar{\bm{\sigma}}^{*},$ (3.6)
    and the method seeks to minimize the discrepancy between the smooth stress $\bm{\sigma}^{*}$ and the stress of the finite element $\bm{\sigma}_{h}$ by enforcing the condition
[…]
    By combining equations (3.6) and (3.7), we can derive the following expression for $\bar{\bm{\sigma}}^{*}$
[…]
    This method is highly accurate but computationally expensive due to the need to solve the additional system of equations.
  • 2.
    The nodal averaging method
    In this simpler approach, the smooth stress $\bar{\sigma}^{*}$ is obtained by averaging the finite element stresses $\bm{\sigma}_{h,i}$ at the nodes
    $\bm{\sigma}^{*}=\frac{\left(\sum_{i=1}^{M}\bm{\sigma}_{h,i}\right)}{M}$ (3.9)
    where $M$ is the number of elements connected with the node $i$.
    This method is straightforward and computationally efficient, although it may not achieve the same level of accuracy as the transformation method. Nevertheless, it performs well in practical applications.
As for error estimation, two types of errors are usually considered. One is the error in $\bm{\sigma}$ which can be expressed as
$\lVert\bm{e}_{\sigma}\rVert=\lVert\bm{\sigma}-\bm{\sigma}_{h}\rVert.$ (3.10)
The other is the recovery-type estimator of the error is similarly defined in the appropriate norm as
$\lVert\tilde{\bm{e}}_{\sigma}\rVert=\lVert\bm{\sigma}^{*}-\bm{\sigma}_{h}\rVert,$ (3.11)
and its efficiency is gauged by the efficacy index defined as
$\theta=\frac{\lVert\tilde{\bm{e}}_{\sigma}\rVert}{\lVert\bm{e}_{\sigma}\rVert}.$ (3.12)
The confines of this efficacy index are elucidated through the subsequent theorem (as detailed in [29]):
$1-\frac{\lVert\bm{e}_{\sigma}^{*}\rVert}{\lVert\bm{e}_{\sigma}\rVert}\leq\theta\leq 1+\frac{\lVert\bm{e}_{\sigma}^{*}\rVert}{\lVert\bm{e}_{\sigma}\rVert},$ (3.13)
with the error in the reconstructed solution characterized as
$\lVert\bm{e}_{\sigma}^{*}\rVert=\lVert\bm{\sigma}-\bm{\sigma}^{*}\rVert.$ (3.14)
It is evident that the efficacy index approaches unity when the precision of the reconstructed solution significantly outperforms that of the original finite element solution.
Should the reconstructed solution exhibit superconvergence, that is,
$\lVert\bm{e}_{\sigma}^{*}\rVert=C^{*}h^{p+\alpha},$ (3.15)
where $\alpha$ is a positive constant and $h$ represents the element size, and
$\lVert\bm{e}_{\sigma}\rVert=Ch^{p},$ (3.16)
then the bounds for the efficacy index are
$1-\tilde{C}h^{\alpha}\leq\theta\leq 1+\tilde{C}h^{\alpha},$ (3.17)
with $\tilde{C}=C^{*}/C$, rendering the error estimator asymptotically precise. This estimation underscores the efficiency of the recovery-type a-posteriori estimator. Given that the error estimation revolves around stress $\bm{\sigma}$, which is contingent upon the gradients of the solution $\bm{u}$, this class of estimator may also be regarded as a form of gradient recovery.
3.2 Algorithm of Recovery-type estimator enhanced PINN
@@
In this section, we introduce the algorithm of recovery-type estimator enhanced adaptive PINN (R-PINN). The success of the adaptive PINNs algorithm hinges on two crucial components: the posterior error estimation method and the adaptive sampling strategy. The posterior error estimation predicts the discrepancy between the numerical and true solutions, while the adaptive sampling strategy adjusts the distribution of collocation points based on this error, allowing for more focused and efficient training in areas that need it most. Section 3.1 provides a comprehensive explanation of the posterior error estimation method. Therefore, for our R-PINN, the key is how to combine PINNs with this kind of posterior error estimator to realize the adaptive process.
To solve the equation (2.2), we utilize a shape-regular triangulation $\mathcal{T}_{h}$ formed by a family of partitions of the domain $\Omega$, which allows us to effectively implement the recovery-type a-posteriori estimator. We define the operator $\mathcal{P}(\hat{u}_{\tilde{\Theta}},\mathcal{T}_{h})$ as the recovery-type estimator used in our algorithm, employing a nodal averaging method for gradient recovery (other gradient recovery methods can also be utilized). The approximated solution $\hat{u}_{\tilde{\Theta}}$ predicts the solution values at all nodes within the mesh $\mathcal{T}_{h}$, which then facilitates gradient recovery and error computation for each element $K\in\mathcal{T}_{h}$ through the recovery-type operator $\mathcal{P}(\hat{u}_{\tilde{\Theta}},\mathcal{T}_{h})$.
Fig. 3: The flow chart of R-PINN.
Another crucial aspect is the selection of collocation points and the adaptive adjustment of their distribution. Similar to many adaptive PINNs algorithms, we start with $N_{1}$ background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$ for pre-training. During the adaptive training phase, we introduce $N_{2}$ adaptive collocation points $\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$ alongside the background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$, updating the adaptive collocation points after each training batch. For background collocation points, we utilize a Sobol sequence as described in section 2.2. For the adaptive collocation points, we propose a novel approach called the recovery-type estimator based adaptive distribution (RE-AD).
Unlike common sampling methods, the RE-AD technique is element-based, as the computed a-posteriori errors are distributed across elements. This means we only need to specify the number of adaptive collocation points to sample in each element, allowing for random sampling within each element. Therefore, it’s vital to allocate adaptive collocation points intelligently to target areas with significant errors without creating redundancy. To achieve this, we design an iterative process: we define a tolerance $\epsilon$ such that $\frac{N_{2}}{N_{e}}\leq\epsilon<1$, where $N_{e}$ is the total number of elements. Initially, we select $L_{1}=[\epsilon N_{e}]$ elements that exhibit the largest errors in $\mathcal{T}_{h}$, which we refer to as the sequence $\mathcal{S}_{1}$. Adaptive collocation points are sampled from $\mathcal{S}_{1}$ in proportion to the errors. For instance, for the $i$-th element in $\mathcal{S}_{1}$, the number of adaptive collocation points assigned is $[N_{2}\frac{e_{i}}{\sum_{j}e_{j}}]$. Since this process rounds down, some adaptive collocation points may remain unallocated. We then repeat this process by selecting $L_{2}=[\epsilon L_{1}]$ of the largest elements and allocating adaptive collocation points according to their errors, continuing until no points are redundant. The RE-AD sampling method differs from RAR but is more akin to RAD, as it maintains a fixed total number of adaptive collocation points while only adjusting their distribution during each iteration.
@@
Algorithm 3 Recovery-type estimator based adaptive distribution (RE-AD)
1:  Set the number of redundant adaptive collocation points $N_{r}=N_{2}$.
2:  Denote $L_{0}=N_{e}$.
3:  $k\leftarrow 1$.
4:  while $N_{r}$ not 00 do
@@
5:     Select $L_{k}=[\epsilon L_{k-1}]$ elements with the largest errors as $\mathcal{S}_{k}$.
6:     Designate adaptive collocation points to $\mathcal{S}_{k}$.
7:     Update the number of redundant adaptive collocation points $N_{r}$.
8:     $k\leftarrow k+1$.
9:  end while
10:  Output adaptive collocation points $\{\bm{x}_{a}\}_{i=1}^{N_{2}}$.
To introduce the specific procedures of the recovery-type estimator enhanced adaptive PINN (R-PINN), we utilize the recovery-type operator $\mathcal{P}(\hat{u}_{\tilde{\Theta}},\mathcal{T}_{h})$ along with the recovery-type estimator based adaptive distribution (RE-AD) sampling strategy. Initially, we use a set of background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$ to train the neural network, aiming to obtain a rough approximation of the true solution, which serves as the pre-training step. Following the pre-training phase, an adaptive training iteration process begins. In each iteration, the recovery-type operator $\mathcal{P}(\hat{u}_{\tilde{\Theta}},\mathcal{T}_{h})$ is used to compute the errors of the approximated solution $\hat{u}_{\tilde{\Theta}}(\bm{x})$ on the mesh $\mathcal{T}_{h}$. Based on these errors, the RE-AD sampling method is then employed to strategically distribute additional adaptive collocation points. These adaptive collocation points are combined with the background collocation points to retrain the neural network. This iterative process continues until the accuracy improvements become minimal, indicating that the network has converged to a satisfactory approximation of the solution. Since the number of collocation points can always stay constant, the algorithm satisfies the characteristics of self-adaptation while ensuring accurate solution.
@@
0:  The number of adaptive samples and background samples $N_{a}$ and $N_{b}$, the maximum steps number $M$
1:  Generate $N_{1}$ background samples $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$ with Sobol sequence.
2:  Pre-train the network $\hat{u}_{\tilde{\bm{\Theta}}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\tilde{\bm{\Theta}})$ with background samples $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$.
3:  Set the number of iteration $k\leftarrow 1$ and give the max number of iterations $M$.
@@
5:     Generate $N_{2}$ adaptive samples $\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$ with RE-AD method.
6:     Retrain the network $\hat{u}_{\tilde{\bm{\Theta}}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\tilde{\bm{\Theta}})$ with the dataset $\mathcal{D}=\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}\cup\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$.
7:  end while
[…]

### v2

@@ 3 Recovery-type estimator enhanced PINN
In this section, we present an adaptive PINN algorithm, which utilizes the recovery-type a-posteriori error estimator [39, 40, 41], a widely used technique in adaptive finite element methods. The concept of adaptive finite element method, first introduced by I. Babuška et al. in 1978 [3], has since seen significant advancements and remains integral to practical scientific computations. This approach extends classical finite element methods by incorporating a-posteriori error estimation and automated mesh refinement to iteratively improve the accuracy of approximated solutions, which can achieve the target accuracy with reduced computational effort, offering distinct advantages over traditional finite element methods, particularly when addressing problems involving singular solutions.
The fundamental concept of the adaptive finite element method lies in the use of a-posteriori error estimator to assess the numerical solution error within each mesh element, thereby guiding the mesh refinement and coarsening process. This adaptive mesh strategy shares a notable similarity with adaptive PINNs, where the distribution of collocation points is adjusted based on the error distribution. In adaptive finite element methods, commonly employed a-posteriori error estimators can be categorized into two main types: the first is based on residual type [3, 4], where the error is estimated by computing the discrepancy between the solution and the residual of the governing equation; the second type, recovery type, proposed by Zienkiewicz and Zhu [39, 40, 41], is a gradient reconstruction-based method that provides more accurate error estimates by reconstructing the solution gradients.
In existing adaptive PINNs algorithms, the PDE residual, computed using automatic differentiation, is typically used as the default a-posteriori error estimator, which corresponds to the residual-based error estimator in adaptive finite element methods. Our proposed algorithm seeks to introduce the recovery-type error estimator into the adaptive PINNs framework, with the aim of enriching and extending the theoretical foundation of adaptive PINNs, ultimately providing a more comprehensive approach to error estimation and control.
@@ 3.1 Recovery-type a-posteriori estimator
In this section, we provide a brief overview of recovery-type a-posteriori error estimators in the context of finite element methods, along with a discussion of how they work in practice. A more detailed theoretical treatment can be found in [36, 31].
Let $\mathcal{T}_{h}$ be a quasi-uniform triangulation of the domain $\Omega$, and denote by $\mathcal{Z}_{h}$ the set of its vertices. For a given node $z\in\mathcal{Z}_{h}$, we define an element patch $\omega$ consisting of all elements that share the vertex $z$. We place the origin of a local coordinate system at $z$, and denote by $(x_{j},y_{j})$ the barycenter of each triangle $\tau_{j}\subset\omega$, where $j=1,2,\dots,m$.
As a model problem, we consider the weak form of a non-self-adjoint problem: Find $u\in H^{1}(\Omega)$ such that
$B(u,v)=\int_{\Omega}\left[(\mathcal{D}\nabla u+\mathbf{b}u)\cdot\nabla v+cuv\right]\,\mathrm{d}{\bm{x}}=f(v),\quad\forall v\in H^{1}(\Omega),$
with a Dirichlet boundary condition $u|_{\partial\Omega}=g$, where $\mathcal{D}$ is a symmetric positive-definite $2\times 2$ matrix, $\mathbf{b}\in[L^{\infty}(\Omega)]^{2}$ is a vector-valued function, $f\in(H^{1}(\Omega))^{\prime}$ is a linear functional, and $g\in H^{\frac{1}{2}}(\partial\Omega)$ is a prescribed function defined on the boundary. We assume that the coefficients are smooth, and that the bilinear form $B(\cdot,\cdot)$ is continuous and satisfies an inf-sup condition on $H^{1}(\Omega)$, ensuring the well-posedness of the problem.
The finite element approximation $u_{h}\in\mathcal{V}_{h}\subset H^{1}(\Omega)$ is defined by
$B(u_{h},v_{h})=f(v_{h}),\quad\forall v_{h}\in\mathcal{V}_{h}.$
To guarantee a unique discrete solution, we further assume that the inf-sup condition also holds on the discrete space $\mathcal{V}_{h}$.
To indicate the recovery-type a-posteriori error, we introduce a gradient recovery operator $G$ designed to provide an improved approximation of the true gradient $\nabla u$. Ideally, the recovered gradient $Gu_{h}$ satisfies the following inequality:
$\lVert\nabla u-Gu_{h}\rVert_{0,\Omega}\leq\beta\lVert\nabla u-\nabla u_{h}\rVert_{0,\Omega},\quad\text{with }0\leq\beta<1,$
which leads to a two-sided bound:
$(1-\beta)\lVert\nabla u-\nabla u_{h}\rVert_{0,\Omega}\leq\lVert\nabla u_{h}-Gu_{h}\rVert_{0,\Omega}\leq(1+\beta)\lVert\nabla u-\nabla u_{h}\rVert_{0,\Omega}.$
This justifies the use of $\lVert\nabla u_{h}-Gu_{h}\rVert_{0,\Omega}$ as a recovery-based a-posteriori error estimator—it is fully computable and reliably reflects the true error.
There are several standard approaches for constructing $Gu_{h}$ at a vertex $z$:
  1. (a)
    Weighted averaging: The recovered gradient is constructed as a weighted average of the gradients evaluated at the barycenters of the neighboring triangles:
    $Gu_{h}(z)=\sum_{j=1}^{m}\frac{|\tau_{j}|}{|\omega|}\nabla u_{h}(x_{j},y_{j}).$
  2. (b)
    Local $L^{2}$ projection: For each component $l=1,2$, we seek a linear function $p_{l}\in P_{1}(\omega)$ satisfying
    $\int_{\omega}\left[p_{l}(x,y)-\partial_{l}u_{h}(x,y)\right]q(x,y)\,\mathrm{d}x\mathrm{d}y=0\quad\forall q\in P_{1}(\omega),$
    and define $Gu_{h}(z)=(p_{1}(0,0),p_{2}(0,0))$.
  3. (c)
    Local least-squares fitting (Zienkiewicz-Zhu method [40]): This method solves a discrete version of the $L^{2}$ projection, by minimizing the difference at the barycenters:
    $\sum_{j=1}^{m}\left[p_{l}(x_{j},y_{j})-\partial_{l}u_{h}(x_{j},y_{j})\right]q(x_{j},y_{j})=0,\quad\forall q\in P_{1}(\omega),\quad l=1,2.$
    Again, we define $Gu_{h}(z)=(p_{1}(0,0),p_{2}(0,0))$.
Subsequently, $Gu_{h}$ can be linearly interpolated from $Gu_{h}(z),\ z\in\mathcal{Z}_{h}$. Among these three approaches, the weighted averaging method is particularly popular due to its simplicity and low computational cost. The least-squares fitting can be viewed as a discrete analogue of the $L^{2}$ projection. Detailed analysis on the convergence and performance of these estimators can be found in [40, 41, 36, 31].
3.2 Algorithm of the recovery-type estimator enhanced PINN
@@
In this section, we introduce the algorithm of recovery-type estimator enhanced adaptive PINN (R-PINN). The success of the adaptive PINNs algorithm hinges on two crucial components: the a-posteriori error estimation method and the adaptive sampling strategy. The a-posteriori error estimation predicts the discrepancy between the numerical and true solutions, while the adaptive sampling strategy adjusts the distribution of collocation points based on this error, allowing for more focused and efficient training on areas that need it the most.
We begin by discussing a-posteriori error estimation based on the recovery-type estimator. The core idea is to utilize the estimator introduced in Section 3.1, particularly the weighted averaging scheme, to assist in evaluating the a-posteriori error. The procedure is as follows. After training the PINN, we obtain an approximated solution $\hat{u}_{\bm{\Theta}}=\mathcal{N}^{L}(\bm{x};\bm{\Theta})$. Given a quasi-uniform triangulation $\mathcal{T}_{h}$ of the domain $\Omega$, we interpolate $\hat{u}_{\bm{\Theta}}$ at mesh vertices to obtain a piecewise linear approximation $\hat{u}_{I}$ such that
$\hat{u}_{I}(z)=\hat{u}_{\bm{\Theta}}(z),\quad\forall z\in\mathcal{Z}_{h},$
where $\mathcal{Z}_{h}$ denotes the set of mesh vertices. The interpolated solution $\hat{u}_{I}$ serves as an intermediate quantity for both gradient recovery and subsequent a-posteriori error estimation. The recovered gradient at a node $z\in\mathcal{Z}_{h}$ is computed using a weighted averaging approach,
$G\hat{u}_{I}(z)=\sum_{j=1}^{m}\frac{|\tau_{j}|}{|\omega|}\nabla\hat{u}_{I}(x_{j},y_{j}),$
where $(x_{j},y_{j})$ is the barycenter of each triangle $\tau_{j}\subset\omega$, and $\omega$ is an element patch consisting of all elements that share the vertex $z$. The recovered gradient field $G\hat{u}_{I}$ is then obtained by linear interpolation of the nodal values $G\hat{u}_{I}(z)$. Using the recovered gradient, the recovery-type a-posteriori error estimator is defined element-wise as
$\eta_{K}=\lVert\nabla\hat{u}_{I}-G\hat{u}_{I}\rVert_{0,K},\quad K\in\mathcal{T}_{h}.$ (3.1)
Here, $\eta_{K}$ provides a local indication of the error within element $K$, guiding the identification of regions where additional sampling points are required to enhance the solution accuracy.
Before presenting the sampling strategy, we outline the overall training strategy. The training process is divided into two phases: a pre-training phase and an adaptive training phase.
In the pre-training phase, the PINN is trained using only a fixed set of background collocation points, denoted by $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$. This phase aims to obtain an initial approximation of the solution. The pre-training is followed by an adaptive training phase, during which both background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$ and a set of adaptively distributed points $\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$ are employed. While the background collocation points remain fixed throughout the entire training process, the adaptively distributed points are dynamically updated in each iteration. Specifically, the adaptive phase comprises $M$ iterations. At the beginning of each iteration, the recovery-based a-posteriori error estimator $\eta_{K}$ is computed to evaluate the current approximation error over each element $K$. Based on these error indicators, new adaptively distributed points are selected in regions with relatively large errors. The PINN is then retrained using the union of background and updated adaptively distributed points for a fixed number of epochs before proceeding to the next iteration.
During the adaptive training process, the sampling strategy plays a key role in determining how the collocation points are selected and updated. The background collocation points, denoted by $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$, are generated using a Sobol sequence as described in Section 2.2, and remain fixed throughout the entire training procedure. In contrast, the adaptively distributed points $\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$ are updated in each iteration based on the element-wise a-posteriori error indicators $\eta_{K}$}. To support the adaptively distributed points selection, we introduce a sampling approach called recovery-type estimator based adaptive distribution (RecAD), which helps allocate adaptively distributed points to regions with larger estimated errors.
Fig. 2: Flowchart of the R-PINN algorithm.
Unlike common sampling methods, the RecAD technique is element-based, as the computed a-posteriori errors are distributed across elements. This means we only need to specify the number of adaptively distributed points to sample in each element, allowing for random sampling within each element. Therefore, a reasonable scheme is required to assign the number of adaptively distributed points across elements based on their estimated a-posteriori errors. To this end, we design an iterative procedure.
We begin by sorting all elements in descending order according to their estimated a-posteriori errors, and denote them by the set $\mathcal{S}_{0}=\{K_{j}\}_{j=1}^{N_{e}}$, where $N_{e}$ is the total number of elements. Let $n_{1},n_{2},\dots,n_{N_{e}}$ denote the number of adaptively distributed points assigned to each element, which are initially all set to zero. Additionally, we introduce a tolerance parameter $\epsilon\in\left(\frac{1}{N_{e}},1\right)$, representing the proportion of elements to be selected for sampling.
@@
Algorithm 1 Recovery-type estimator based adaptive distribution (RecAD)
1: Initialize the number of remaining adaptively distributed points $N_{r}=N_{2}$, the number of remaining elements $m_{0}=N_{e}$, and set the tolerance parameter $\epsilon$.
2: Sort all elements in descending order according to their estimated a-posteriori errors to obtain the set $\mathcal{S}_{0}=\{K_{j}\}_{j=1}^{N_{e}}$.
3: Set the initial number of assigned points for each element as $n_{j}=0$, for $j=1,2,\cdots,N_{e}$.
4: Set iteration counter $k\leftarrow 1$.
5:while $N_{r}\neq 0$ do
@@
6:  Select $m_{k}=\left[\epsilon m_{k-1}\right]$ elements with the largest a-posteriori errors from $\mathcal{S}_{k-1}$ as $\mathcal{S}_{k}$.
7:  Assign the number of adaptively distributed points to elements in $\mathcal{S}_{k}$ proportionally:
$n_{j}\leftarrow n_{j}+\left[N_{r}\frac{\eta_{K_{j}}}{\sum_{i=1}^{m_{k}}\eta_{K_{i}}}\right],\quad j=1,2,\cdots,m_{k}.$
8:  Update the number of remaining adaptively distributed points $N_{r}$ by equation (3.2).
9:  if $m_{k}<\frac{1}{\epsilon}$ then
10:   Assign the remaining points to the element with the largest error: $n_{1}\leftarrow n_{1}+N_{r}$.
11:   break
12:  end if
13:  $k\leftarrow k+1$.
14:end while
15: Randomly sample the adaptively distributed points $\{\bm{x}_{a}\}_{i=1}^{N_{2}}$ with the numbers $n_{1},n_{2},\cdots,n_{N_{e}}$ in each element.
16:Output: Adaptively distributed points $\{\bm{x}_{a}\}_{i=1}^{N_{2}}$.
At the first iteration, we select the top $m_{1}=\left[\epsilon N_{e}\right]$ elements with the largest a-posteriori errors from $\mathcal{S}_{0}$, and denote this subset as $\mathcal{S}_{1}=\{K_{j}\}_{j=1}^{m_{1}}$. In these selected elements, the number of adaptively distributed points are assigned approximately in proportion to their estimated a-posteriori errors,
$n_{j}=\left[N_{2}\frac{\eta_{K_{j}}}{\sum_{i=1}^{m_{1}}\eta_{K_{i}}}\right],\quad j=1,2,\cdots,m_{1}.$
Since rounding is applied to ensure that each $n_{j}$ is an integer, the total number of assigned points is likely to be smaller than the prescribed $N_{2}$. We then define the number of remaining points $N_{r}$ as
$N_{r}=N_{2}-\sum_{j=1}^{N_{e}}n_{j}.$ (3.2)
In the next iteration, we select a subset of elements from $\mathcal{S}_{1}$, specifically the top $m_{2}=\left[\epsilon m_{1}\right]$ elements with the largest a-posteriori errors, denoted as $\mathcal{S}_{2}=\{K_{j}\}_{j=1}^{m_{2}}$. The number of remaining points $N_{r}$ are then redistributed to these elements according to their relative a-posteriori errors
$n_{j}=n_{j}+\left[N_{r}\frac{\eta_{K_{j}}}{\sum_{i=1}^{m_{2}}\eta_{K_{i}}}\right],\quad j=1,2,\cdots,m_{2}.$
After each iteration, we update $N_{r}$ using the same equation (3.2). The process continues until one of the following stopping criteria is met: (1) the number of adaptively distributed points $N_{2}$ has been fully allocated (namely $N_{r}=0$), or (2) the number of selected elements becomes smaller than $\frac{1}{\epsilon}$, meaning the next subset would be empty. In the second case, suppose that after $k$-th iterations we obtain $\mathcal{S}_{k}=\{K_{j}\}_{j=1}^{m_{k}}$ with $N_{r}>0$ and $m_{k}<\frac{1}{\epsilon}$. Then, continuing the iteration would result in an empty set $\mathcal{S}_{k+1}$. To avoid this, we directly assign the number of remaining points $N_{r}$ to the element with the largest a-posteriori error $K_{1}$, namely $n_{1}=n_{1}+N_{r}$.Now that the number of adaptively distributed points for each element in the mesh $\mathcal{T}_{h}$ has been fully determined, we then randomly sample the corresponding number of points within each element based on the assigned values $n_{1},n_{2},\dots,n_{N_{e}}$, thereby completing the adaptive sampling process.
The full algorithm for this adaptive distribution scheme RecAD is summarized in Algorithm 1. With RecAD in place, all components of the R-PINN framework are complete. The entire training procedure of R-PINN is outlined in Algorithm 2.
@@
0: Number of adaptively distributed points $N_{a}$, number of background points $N_{b}$, and maximum number of adaptive iterations $M$.
1: Generate $N_{1}$ background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$ with Sobol sequence.
2: Pre-train the network $\hat{u}_{\bm{\Theta}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\bm{\Theta})$ with background collocation points $\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}$.
3: Set the iteration counter $k\leftarrow 1$.
@@
5:  Compute the a-posteriori errors of $\hat{u}_{\bm{\Theta}}(\bm{x})$ with the recovery-type estimator defined in the equation (3.1).
6:  Generate $N_{2}$ adaptively distributed points $\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$ with RecAD method.
7:  Retrain the network $\hat{u}_{\bm{\Theta}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\bm{\Theta})$ with the dataset $\mathcal{D}=\{\bm{x}_{b}^{(i)}\}_{i=1}^{N_{1}}\cup\{\bm{x}_{a}^{(i)}\}_{i=1}^{N_{2}}$.
8:end while
9:return The approximated solution $\hat{u}_{\bm{\Theta}}(\bm{x})=\mathcal{N}^{L}(\bm{x};\bm{\Theta})$.